OpenStudy (anonymous):

Medal included and fan!! Solve please. 5x + 10 = −4x − 17 x =

OpenStudy (ageta):

x=−3

OpenStudy (ageta):

Let's solve your equation step-by-step. 5x+10=−4x−17 Step 1: Add 4x to both sides. 5x+10+4x=−4x−17+4x 9x+10=−17 Step 2: Subtract 10 from both sides. 9x+10−10=−17−10 9x=−27 Step 3: Divide both sides by 9.

OpenStudy (ageta):

the answer is x-3

OpenStudy (anonymous):

5x + 10 = -4x - 17 - 10 = - 10 5x = -4x - 27 +4x = + 4x 9x = -27 /9 = /9 x = -3

OpenStudy (anonymous):

thank you!!

OpenStudy (ageta):

your welcome